I need help with dividing fractions, I am very confused on how to simplify my problem and keep it in the indicated operation. Here is an example of what I mean;
4/3 divided by 5/7?
Please get back to me asap
Thanks alot!

Answer by stanbon(75887)

4/3 divided by 5/7?
---
Multiply numerator and denominator by (7/5) to get:
---
[(7/5)(4/3)]/[(7/5)(5/7)]
----
= (28/15)/1
---
= (28/15)
---------------------
Note: The general rule for dividing a fraction
by a fraction is "invert the denominator and multiply".
-----
Example: (2/3)/(5/4) = (2/3)(4/5) = 8/15
===============================================
Cheers,
Stan H.